# Vision-LSTM: xLSTM as Generic Vision Backbone

### v2

@@ Abstract
Transformers are widely used as generic backbones in computer vision, despite initially introduced for natural language processing. Recently, the Long Short-Term Memory (LSTM) has been extended to a scalable and performant architecture – the xLSTM – which overcomes long-standing LSTM limitations via exponential gating and parallelizable matrix memory structure. In this report, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision. ViL comprises a stack of xLSTM blocks where odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. Experiments show that ViL holds promise to be further deployed as new generic backbone for computer vision architectures.
@@ 2 Method
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1. Following ViT [18], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Figure 1: Schematic overview of Vision-LSTM (ViL). Following ViT [18], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL uses the concatenation of the first and the last patch as input to a linear classification head.
@@ 3 Experiments
As ViL is an isotropic architecture (like ViT [18]), we compare it against other isotropic architectures and leave exploration of hierarchical ViL architectures to future work.
Figure 2: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute. ViL shows strong performances across classification and semantic segmentation tasks.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [18, 51, 53, 52, 54, 27]. Since this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases. Even so, Figure 2 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim) [67] on ImageNet-1K [17] classification, ADE20K [66] semantic segmentation and VTAB-1K [64] classification. Detailed results are presented in the following sections.
3.1 ImageNet-1K Pre-training
Table 1 relates parameter counts and FLOPS to classification accuracy after pre-training on ImageNet-1K. ViL outperforms heavily optimized ViT protocols and other backbones on the tiny and small scale. While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 3 and Table 2) show that ViL-B still learns strong features. Additionally, training a ViL-B for 400 epochs takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs. Therefore, the hyperparameter configuration is most likely not optimal as we were not able to extensively optimize hyperparameters for this scale. Note that training ViL is still quite expensive to train as there does not exist an optimized hardware implementation yet, which does exist for other models like ViT [15, 14] or Vim [23, 67]. Nevertheless, the efficient alternating block design of ViL makes it already faster than Vim (up to 69% as shown in Appendix B.1) while only using torch.compile, a generic speed optimization of the PyTorch [45] framework.
@@ 3.1 ImageNet-1K Pre-training
Mamba®-B [58] 280+20 99M - 82.9
ViL-B 400+5 89M 18.6G 82.4
Table 1: ImageNet-1K pre-training accuracy. All models use a patch size of 16x16 with 224x224 resolution at most. Models with “+” in their “Epochs” column pre-train on lower resolution followed by fine-tuning on 224x224 resolution for some epochs. ViL performs favorably against against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT), RWKV [47] (VRWKV) and Mamba [23] (Vim, Mamba®).
@@ 3.2 ADE20K Semantic Segmentation
Table 2 shows results for transferring ImageNet-1K pre-trained models to ADE20K [66] semantic segmentation using UperNet [61]. Also here, ViL shows strong performances across the board, even outperforming DeiT-III-B despite the lower ImageNet-1K accuracy of ViL-B.
For this benchmark, we train on 512x512 images and inference could be conducted on the original images (without resizing), which results in even larger images. Consequently, the number of patch tokens can become quite large. As sequence length increases, ViL would ideally use a recurrent or chunkwise formulation of the mLSTM layer. Shortly summarized, the mLSTM can operate in parallel, recurrent or chunked modes, each with distinct FLOPS and runtime characteristics. Given a sequence length $L$ and hidden dimension $D$, the trade-off between modes can be formulated as $\mathcal{O}(LCD+LD^{2})$ where $C$ denotes the chunksize of the chunked form. $C=1$ (recurrent mode) yields complexity $\mathcal{O}(LD^{2})$ and $C=L$ (parallel mode) gives $\mathcal{O}(L^{2}D)$ or $\mathcal{O}(L\frac{L}{2}D)$ when omitting the causally masked operations. Due to this trade-off and the current lack of an optimized hardware implementation for the mLSTM, it is difficult to create runtime or FLOPS estimates that are of practical relevance in this setting and therefore leave it to future work.
@@
Model #Params mIoU ACC mIoU ACC
DeiT-T [51] 10M 38.1 78.2 40.3 79.9
DeiT-III-T [54] 10M 39.8 79.2 42.2 80.7
Vim-T [67] 13M 41.0 - - -
ViL-T 11M 41.2 80.2 43.1 81.3
DeiT-S [51] 41M 43.1 80.7 45.2 81.8
DeiT-III-S [54] 41M 45.2 81.5 46.3 82.3
Vim-S [67] 46M 44.9 - - -
Mamba®-S [58] 56M 45.3 - - -
ViL-S 42M 46.3 82.0 47.9 82.9
DeiT-B [51] 113M 45.8 82.1 47.0 82.9
DeiT-III-B [54] 113M 47.5 82.6 49.0 83.3
Mamba®-B [58] 132M 47.7 - - -
ViL-B 115M 48.6 82.8 49.6 83.3
Table 2: Semantic segmentation results on ADE20K [66] using UperNet [61]. We report mean intersection over union (mIoU) and pixelwise accuracy (ACC) for single- and multi-scale evaluation. Models are trained for 160K updates with a batchsize of 16 on 512x512 resolution. Detailed hyperparameters are listed in Appendix Table 14.
@@ 3.3 VTAB-1K Transfer Classification
Table 3 shows transfer classification results for ImageNet-1K pre-trained models on the VTAB-1K [64] benchmark. VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [38], Caltech101 [20] or SUN397 [60]), 4 specialized datasets (medical imaging [57, 34] and remote sensing [28, 11]) and 8 structured datasets (with tasks such as object counting [33] or binned depth estimation [21]). We follow common practices and tune the learning rate per model and dataset on the validation set followed by training with the best learning rate on the union of train and validation set. The performance metric is the average testset accuracy over 5 seeds. ViL shows strong transfer classification performance outperforming all other models on the average over all 19 datasets. ViL performs particularly well on the structured datasets where ViL-B outperforms DeiT-III-B despite ViL-B having lower ImageNet-1K accuracy.
@@
DeiT-T [51] 6M 1.3G 69.2 82.0 53.3 65.2
DeiT-III-T (reimpl.) 6M 1.3G 71.9 82.6 55.2 67.1
Vim-T [67] 7M 1.5G 68.0 80.7 47.1 61.9
ViL-T 6M 1.5G 73.6 83.4 56.1 68.3
DeiT-S [51] 22M 4.6G 73.3 83.8 53.2 67.1
DeiT-III-S [54] 22M 4.6G 75.0 83.2 52.3 67.2
Vim-S [67] 26M 5.3G 69.6 81.7 49.4 63.6
ViL-S 23M 5.1G 75.3 84.3 58.3 70.0
DeiT-B [51] 86M 17.6G 76.5 85.2 55.7 69.6
DeiT-III-B [54] 86M 17.6G 77.6 84.8 56.6 70.3
ViL-B 89M 18.6G 76.6 84.7 59.1 70.9
@@
Table 3: Transfer classification accuracies on the VTAB-1K [64] benchmark using ImageNet-1K pre-trained models. VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets. We show averages per category and the average accuracy over all 19 datasets (Appendix Table 10 lists all individual accuracies). ViL shows strong transfer classification performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark. Notably, ViL performs exceptionally well on the structured category. We tune the learning rate for each model and dataset on the validation set and report the average performance on the test set over 5 seeds. Appendix Table 13 lists further hyperparameters.
3.4 Block Design
We study different ways to design ViL blocks in Table 4. The plain, uni-directional xLSTM block fails to reach competitive performances since the autoregressive nature of xLSTM is not suited for image classification. Traversing blocks in a bi-directional manner, i.e., introducing a second mLSTM layer in each block that traverses the sequence backwards (akin to Vim [67]), improves performance, but also requires more parameters and FLOPS. Sharing the parameters of the forward and backward mLSTM makes the model more parameter efficient, but still requires more compute and overloads these parameters which leads to performance drops. Using alternating blocks enhances performance while remaining compute and parameter efficient. We also explore quad-directional designs (similar to [40]), which refers to traversing the sequence row-wise (in both directions) and additionally column-wise (in both directions). Bi-directional traverses the sequence only row-wise (in both directions). Figure 3 visualizes the different traversal paths.
Due to the increased cost of bi-directional and quad-directional blocks, this study was conducted in a heavily reduced setting. We train on a subset of ImageNet-1K that contains only samples from 100 classes for 400 epochs at 128x128 resolution. This was particularly necessary because our quad-directional implementation is not compatible with torch.compile (a generic speed optimization method from PyTorch [45]), which leads to longer runtimes, as indicated last column of Table 4. Due to this technical limitation, we choose the alternating bi-directional block as our core design.
Direction Alternating Param. Sharing ACC #Param FLOPS Runtime
Uni-dir. (xLSTM) 72.8 6.1M 1.0x 1.0x
Bi-dir. 74.8 6.5M 1.4x 1.9x
Bi-dir. 73.3 6.1M 1.4x 1.9x
Bi-dir. (ViL) 74.8 6.1M 1.0x 1.0x
Quad-dir. 76.3 7.3M 2.3x 7.5x
Quad-dir. 72.8 6.1M 2.3x 7.5x
Quad-dir. 75.6 6.1M 1.0x 1.9x
Table 4: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block. However, using multiple directions per block also requires more compute and parameters. Sharing parameters between directions matches #Params but overloads them, leading to performance drops. Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances. While quad-directional designs yields the best performances, they comes with drastically increased runtime. As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions. Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128. “Param. Sharing” shares the parameters of the two (bidirectional) or four (quaddirectional) mLSTM layers in each block.
Figure 3: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.

### v3

@@ Abstract
Transformers are widely used as generic backbones in computer vision, despite initially introduced for natural language processing. Recently, the Long Short-Term Memory (LSTM) has been extended to a scalable and performant architecture – the xLSTM – which overcomes long-standing LSTM limitations via exponential gating and parallelizable matrix memory structure. In this paper, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision. ViL comprises a stack of xLSTM blocks where odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. ViL achieves strong performances on classification, transfer learning and segmentation tasks as well as a beneficial pre-training cost-to-performance trade-off. Experiments show that ViL holds promise to be further deployed as new generic backbone for computer vision architectures.
@@
Figure 1: The efficient and scalable design of Vision-LSTM shows strong performances, uses less FLOPS than Transformer/Mamba counterparts and scales linear to higher resolutions. Performance is averaged over ImageNet accuracy, ADE20K mIoU and VTAB-1K accuracy.
footnotetext: Published as a conference paper at ICLR 2025
@@ 1 Introduction
Figure 2: Schematic overview of Vision-LSTM (ViL). Following ViT [22], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL uses the concatenation of the first and the last patch as input to a linear classification head. ViL is an isotropic architecture, i.e., all blocks have the same input and output dimension and no downsampling layers are used except the initial patch embedding. Projection layers process each patch individually and the mLSTM exchanges information between patches.
@@
Our contributions summarize as follows:
  • We introduce Vision-LSTM (ViL), an adaption of the mLSTM to computer vision tasks that can serve as a generic vision backbone with linear complexity.
  • We show modeling capacity and generalization in the common vision benchmark of pre-training models on ImageNet-1K, followed by fine-tuning on transfer classification and semantic segmentation tasks.
  • We ablate various architectural design choices to evaluate their impact on performance and provide insights into the model design.
  • We discuss potential future directions and current limitations that, once addressed, will improve ViL even further.
@@ 2 Method
Vision-LSTM (ViL) introduces xLSTM [5] to computer vision, similar to other vision adaptions of sequence modeling architectures, e.g., Vision Transformers [22], Vision Mamba [82], or Vision RWKV [23].
2.1 Preliminaries
[…]
The key innovations of the mLSTM [5] are the enhanced storage capacity compared to the classical LSTM [39] by using a matrix memory cell ${\bm{C}}\in\mathbb{R}^{d\times d}$ instead of a scalar memory cell $c\in\mathbb{R}$ and introducing exponential gates (instead of sigmoid gates) to the input and forget gates, where $d$ is the hidden dimension within the mLSTM block (typically $d=2D$).
Intuitively, the mLSTM is a more expressive and faster version of the classical LSTM that can be efficiently parallelized on modern hardware. In ViL, the mLSTM is used to process dependencies between patches, similar to how the attention exchanges information between patches in a ViT. The mLSTM is embedded into a gated MLP architecture, as shown on the right of Figure 2, where the weight matrices of the MLP process each patch individually and the mLSTM exchanges information between patches. For completeness, we outline the forward pass of the mLSTM in the following paragraphs.
The mLSTM [5] is a recurrent neural network, which maps a state $({\bm{h}}_{t-1},{\bm{C}}_{t-1},{\bm{n}}_{t-1})$ to a successor state $({\bm{h}}_{t},{\bm{C}}_{t},{\bm{n}}_{t})$ given input ${\bm{x}}_{t-1}$. Thereby, ${\bm{h}}_{t}\in\mathbb{R}^{d}$ denotes the hidden state, ${\bm{C}}_{t}\in\mathbb{R}^{d\times d}$ is the cell state and ${\bm{n}}_{t}\in\mathbb{R}^{d}$ corresponds to a normalizer state. The full forward pass of the mLSTM is as follows [5]:
${\bm{C}}_{t}$ $=\ f_{t}\ {\bm{C}}_{t-1}\ +\ i_{t}\ {\bm{v}}_{t}\ {\bm{k}}_{t}^{\top}$ cell state (1)
${\bm{n}}_{t}$ $=\ f_{t}\ {\bm{n}}_{t-1}\ +\ i_{t}\ {\bm{k}}_{t}$ normalizer state (2)
${\bm{h}}_{t}$ $=\ {\bm{o}}_{t}\ \odot\ \tilde{{\bm{h}}}_{t}$ $\tilde{{\bm{h}}}_{t}$ $=\ {\bm{C}}_{t}{\bm{q}}_{t}\ /\ \max\left\{|{{\bm{n}}_{t}^{\top}{\bm{q}}_{t}}|,1\right\}$ hidden state (3)
${\bm{q}}_{t}$ $=\ {\bm{W}}_{q}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{q}$ query input (4)
${\bm{k}}_{t}$ $=\ \frac{1}{\sqrt{d}}{\bm{W}}_{k}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{k}$ key input (5)
${\bm{v}}_{t}$ $=\ {\bm{W}}_{v}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{v}$ value input (6)
$i_{t}$ $=\ \exp\!\big{(}\tilde{i}_{t}\big{)}$ $\tilde{i}_{t}$ $=\ {\bm{w}}^{\top}_{i}\ {\bm{x}}_{t}\ +\ b_{i}$ input gate (7)
$f_{t}$ $=\ \exp\!\big{(}\tilde{f}_{t}\big{)}$ $\tilde{f}_{t}$ $=\ {\bm{w}}^{\top}_{f}\ {\bm{x}}_{t}\ +\ b_{f}$ forget gate (8)
${\bm{o}}_{t}$ $=\ \sigma\big{(}\tilde{fo}_{t}\big{)}$ $\tilde{{\bm{o}}}_{t}$ $=\ {\bm{W}}_{{\bm{o}}}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{{\bm{o}}}$ output gate (9)
As exponential activation functions can lead to large activations, the input and forget gates are stabilized with an additional state $m_{t}$:
$m_{t}$ $=\max\!\Big{(}\log(f_{t})+{\bm{m}}_{t-1},\log(f_{t})\Big{)}$ stabilizer state (10)
$i^{\prime}_{t}$ $=\exp\!\Big{(}\log(i_{t})-m_{t}\Big{)}=\exp\Big{(}\tilde{i}-m_{t}\Big{)}$ stabilized input gate (11)
$f^{\prime}_{t}$ $=\exp\!\Big{(}\log(f_{t})+m_{t-1}-m_{t}\Big{)}$ stabilized forget gate (12)
As the mLSTM has no memory mixing, i.e, interactions between hidden states from one timestep to the next, it can be fully parallelized for fast computation on modern hardware. For a detailed discussion and theory of the cell state update, further details to the mLSTM we refer to the original work [5].
2.2 Vision-LSTM (ViL)
@@
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from mLSTM blocks, as visualized in Figure 2. Following ViT [22], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Formally, the forward pass of a pair of ViL blocks is:
${\bm{Y}}^{\prime}$ $={\bm{X}}+\text{Block}_{\theta}({\bm{X}})$ (13)
${\bm{Y}}$ $={\bm{Y}}^{\prime}+\text{Flip}(\text{Block}_{\phi}(\text{Flip}({\bm{Y}}^{\prime})))$ (14)
Where “Flip” reverses the sequence and “$\text{Block}_{\theta}$” and “$\text{Block}_{\phi}$” corresponds to mLSTM blocks with parameters $\theta$ and $\phi$ (shown in Figure 2, right).
A key motivation of ViL is that the autoregressive mLSTM can operate in a recurrent, parallel or chunkwise mode, each with distinct FLOPS and runtime characteristics. Given a sequence length $T$ and hidden dimension $d$, the complexity of the recurrent mode is $\mathcal{O}(Td^{2})$ and needs to be processed sequentially, whereas the parallel mode has complexity $\mathcal{O}(T^{2}d)$ and is fully parallelizable. The chunkwise mode combines the advantages of the other modes by introducing a chunksize $S$ where the parallel mode is used within chunks and the recurrent mode between chunks. This allows high parallelization, minimal operations and linear scaling with $T$. Complexity wise, the chunkwise mode has $\mathcal{O}(\frac{T}{S}S^{2}d+\frac{T}{S}d^{2})$ or $\mathcal{O}(TSd+\frac{T}{S}d^{2})$ where $\frac{T}{S}$ corresponds to the number of chunks.
@@ 3 Experiments
Figure 3: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute. ViL shows strong performances across classification (ImageNet-1K), semantic segmentation (ADE20K) and transfer classification (VTAB-1K) tasks.
@@
We then transfer the pre-trained models to serveral benchmark tasks: ImageNet-1K classification on the validation set, ADE20K [81] semantic segmentation and VTAB-1K [79] classification. These benchmarks evaluate global image understanding (ImageNet-1K), semantic local and global understanding (ADE20K) and few-shot generalization to a diverse set of 19 VTAB-1K classification datasets, which include natural images, specialized imagery (medical and satellite) and structured tasks (camera angle prediction, depth estimation, object counting, …).
Figure 3 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim). Detailed results are presented in the following sections.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [22, 64, 66, 65, 67]. Therefore, a vast part of the hyperparameter space for pre-training ViTs has been explored. Since this work is the first to apply xLSTM to computer vision, considerably less effort has been put into hyperparameter tuning and architecture optimization, suggesting that future work could improve ViL even further.
3.1 ImageNet-1K Classification
Table 1 relates parameter counts and FLOPS to validation accuracy after pre-training on ImageNet-1K. ViL outperforms heavily optimized ViT protocols and other backbones on the tiny and small scale. While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 2 and Table 3) show that ViL-B still learns strong features, particularly for semantic segmentation and structured tasks.
@@ 3.1 ImageNet-1K Classification
Mamba®-B [73] 280+20 99M 20.6G 82.9
ViL-B 400+5 89M 17.9G 82.4
Table 1: ImageNet-1K pre-training accuracy. All models use a patch size of 16x16 with 224x224 resolution at most. Models with “+” in their “Epochs” column pre-train on lower resolution followed by fine-tuning on 224x224 resolution for some epochs. ViL performs favorably against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT series), RWKV (VRWKV) and Mamba (Vim, Mamba®). Appendix Table 9 confirms these results on OOD and robustness evaluations of these classifiers.
@@ 3.2 ADE20K Semantic Segmentation
Table 2 shows results for transferring ImageNet-1K pre-trained models to ADE20K [81] semantic segmentation using UperNet [75]. Also here, ViL shows strong performances across the board, even outperforming DeiT-III-B despite the lower ImageNet-1K accuracy of ViL-B. The high resolution of the ADE20K segmentation task (512x512) results in a total of 1024 patch tokens where the quadratic complexity of self-attention is significantly more expensive than the linear complexity of the mLSTM, resulting in much fewer FLOPS for ViL. Additionally, the efficient alternating block design results in lower FLOPS than Mamba-based vision models (which also have linear complexity).
@@
Model #Params FLOPS mIoU ACC mIoU ACC
DeiT-T 10M 10.4G 38.1 78.2 40.3 79.9
DeiT-III-T 10M 10.4G 39.8 79.2 42.2 80.7
Vim-T 13M 7.7G 41.0 - - -
ViL-T 11M 6.6G 41.2 80.2 43.1 81.3
DeiT-S 41M 31.7G 43.1 80.7 45.2 81.8
DeiT-III-S 41M 31.7G 45.2 81.5 46.3 82.3
Vim-S 46M 27.3G 44.9 - - -
Mamba®-S 56M 27.6G 45.3 - - -
ViL-S 42M 24.4G 46.3 82.0 47.9 82.9
DeiT-B 113M 107.0G 45.8 82.1 47.0 82.9
DeiT-III-B 113M 107.0G 47.5 82.6 49.0 83.3
Mamba®-B 132M 102.8G 47.7 - - -
ViL-B 115M 93.6G 48.6 82.8 49.6 83.3
Table 2: Semantic segmentation results on ADE20K [81] using UperNet [75]. We report mean intersection over union (mIoU) and pixelwise accuracy (ACC) for single- and multi-scale evaluation. Models are trained for 160K updates with a batchsize of 16 on 512x512 resolution. We use a feature pyramid consisting of rescaled feature maps after the 4th, 6th, 8th and final block. Detailed hyperparameters are listed in Appendix Table 12. FLOPS are calculated only from the backbone at 512x512 resolution as all models use the same segmentation head.
@@ 3.3 VTAB-1K Transfer Classification
DeiT-T 6M 1.3G 69.2 82.0 53.3 65.2
DeiT-III-T 6M 1.3G 71.9 82.6 55.2 67.1
Vim-T 7M 1.5G 68.0 80.7 47.1 61.9
ViL-T 6M 1.3G 73.6 83.4 56.1 68.3
DeiT-S 22M 4.6G 73.3 83.8 53.2 67.1
DeiT-III-S 22M 4.6G 75.0 83.2 52.3 67.2
Vim-S 26M 5.3G 69.6 81.7 49.4 63.6
ViL-S 23M 4.7G 75.3 84.3 58.3 70.0
DeiT-B 86M 17.6G 76.5 85.2 55.7 69.6
DeiT-III-B 86M 17.6G 77.6 84.8 56.6 70.3
ViL-B 89M 17.9G 76.6 84.7 59.1 70.9
@@
Table 3: Transfer classification accuracies on the VTAB-1K [79] benchmark using ImageNet-1K pre-trained models. VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets. We show averages per category and the average accuracy over all 19 datasets (Appendix Table 8 lists all individual accuracies). ViL shows strong generalization performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark. ViL performs exceptionally well on the structured category. We tune the learning rate for each model and dataset on the validation set and report the average testset accuracy over 5 seeds. Appendix Table 11 lists further hyperparameters.
Table 3 shows transfer classification results for ImageNet-1K pre-trained models on the VTAB-1K [79] benchmark. VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [45] or Caltech101 [24]), 4 specialized datasets (medical imaging [71, 43] and remote sensing [35, 14]) and 8 structured datasets (with tasks such as object counting [42] or binned depth estimation [26]). We follow common practices and tune the learning rate per model and dataset on the validation set followed by training with the best learning rate on the union of train and validation set. The performance metric is the average testset accuracy over 5 seeds. ViL shows strong transfer classification performance outperforming all other models on the average over all 19 datasets. ViL performs particularly well on the structured datasets where ViL-B outperforms DeiT-III-B despite ViL-B having lower ImageNet-1K accuracy.
4 Ablation Studies
We ablate various design choices of ViL by training ViL-T models for 100 epochs on ImageNet-1K in 224x224 resolution, other hyperparameters follow the ones from Section 3 (see also Appendix B.3). We then report the validation accuracy on ImageNet-1K and fine-tune the model on ADE20K to ensure that design choices are not overfitted to classification. We also use a reduced segmentation pipeline where we use a linear segmentation head and train for 40K updates using a batch size of 16 (other hyperparameters follow Appendix 12).
4.1 Architectural Design
We consider various architecture design choices in Table 4.
Directions IN1K ADE20K
Uni-dir. 72.2 28.6
Bi-dir. 73.7 31.7
Quad-dir. 73.8 33.1
Oct-dir. 73.5 32.4
Convolution IN1K ADE20K
None 72.3 29.2
Causal-Conv1D 72.8 27.8
Conv1D 72.8 28.4
Conv2D 73.7 31.7
Pos. Embed. IN1K ADE20K
73.7 31.0
73.7 31.7
Concurrency IN1K ADE20K
Sequential 73.7 31.7
Parallel 73.0 30.6
Table 4: Architecture design ablation studies. Default settings
Figure 4: Uni-directional, bi-directional, quad-directional and oct-directional traversal paths. Squares represent individual patch tokens. Traversal starts at the circle and goes in direction of the arrow, if no further patches are in a row/column, the traversal continues in the next row/column as indicated by the dashed line.
(a) Traversal Directions
Traversing the sequence in at least two directions greatly improves performance due to the non-causal 2D structure of images. Adding column-wise traversal directions (Quad-dir.) could even further improve semantic segmentation performance. Additionally using 4 instead of 2 starting positions (Oct-dir.) shows no benefit. Note that all variants have the same amount of FLOPS due to sequential application of different directions. Directions are visualized in Figure 4.
We use “Bi-dir.” for our final models due to current technical limitations which would slow down training on more than 2 directions. This limitation comes from the current lack of optimized hardware implementations of the mLSTM (e.g., CUDA kernels) where we instead rely on torch.compile, a generic speed optimization method from PyTorch [56], to optimize computations. Our implementation of quad- and oct-directional traversals is not compatible with torch.compile, which results in approximately double the runtime. We therefore train all models from Section 3 with “Bi-dir.”. Note that this is merely a technical limitation, not a methodical one and the ablation study suggest that future ViL models could be even better using a quad-directional design.
(b) QK Convolution
The mLSTM block design uses a causal 1D convolution to aggregate local context to improve storage/retrieval to/from the cell state ${\bm{C}}$. This is done by applying a convolution layer to ${\bm{X}}$ before projecting it to ${\bm{Q}}$ with ${\bm{W}}_{q}$ and ${\bm{K}}$ with ${\bm{W}}_{k}$ respectively. The convolution is shared for ${\bm{Q}}$ and ${\bm{K}}$. The causal 1D structure of the convolution from the original mLSTM [5] is necessary due to the causal 1D structure of language modeling. However, as images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3). This allows the mLSTM to make better storage/retrieval decisions through the added local context.
(c) Positional Embedding
ViTs require positional embedding to tell the model where each patch is located in the image, suffering heavy performance losses if the position is not required [22, 15]. The mLSTM is an autoregressive model, which makes it optional to add positional embeddings as it can recognize the position of the current patch based on how many patches have been processed. However, the ablation shows that it is nevertheless beneficial to provide this information explicitly as it improves segmentation results without hurting classification performance.
(d) Sequential vs. Parallel
Related architectures use a parallel design where a sequence is processed from multiple directions in a single block [82, 23]. We investigate a similar design where we apply both directions in parallel instead of sequentially. To keep parameters and FLOPS constant, we apply the directions akin to parallel transformer blocks [72] while halving the depth.
${\bm{Y}}={\bm{X}}+\text{Block}_{\theta}({\bm{X}})+\text{Flip}(\text{Block}_{\phi}(\text{Flip}({\bm{X}})))$ (15)